# Modified scattering for the three dimensional Maxwell-Dirac system

### v4

@@ 4.3. Bounds for derivatives
Using Gagliardo-Nirenberg estimates, we can derive energy estimates for $(\partial^{I}\psi,\partial^{I}A)$ when $I_{0}=0$. The general case can then be derived inductively from the equations in (4.5).
@@
(4.13) $T_{1}^{\frac{1}{2}}\|F_{I}\|_{L^{2}(C_{T_{1}})}+T_{1}^{\frac{1}{3}}\|G_{I}\|_{L^{\frac{3}{2}}(C_{T_{1}})}\lesssim C_{0}\varepsilon\|\partial^{\leq k}(\psi,A)\|_{X_{T_{1}}},\qquad|I|\leq k.$
@@
$T_{1}^{\frac{1}{2}}\|F_{I}\|_{L^{2}(C_{T_{1}})}+T_{1}^{\frac{1}{3}}\|G_{I}\|_{L^{\frac{3}{2}}(C_{T_{1}})}\lesssim C_{0}\varepsilon\|\partial^{\leq k}(\psi,A)(T_{1})\|_{\mathscr{H}^{0}}.$
@@
This last inequality forms the basis for our discrete Gronwall, and, iterating, implies that
$\|\partial^{\leq k}(\psi,A)(t)\|_{\mathscr{H}^{0}}\lesssim\langle t\rangle^{c\varepsilon}\|\partial^{\leq k}(\psi,A)(0)\|_{\mathscr{H}^{0}},\qquad 0\leq t\leq T.$
Together with (4.14), this implies the desired estimate (4.12).
@@
(4.16) $\|F_{I}\|_{L^{2}(C_{T_{1}})}\lesssim\|A\|_{L^{\infty}(C_{T_{1}})}\|\partial^{\leq k}\psi\|_{L^{2}(C_{T_{1}})}+\|\psi\|_{L^{6}(C_{T_{1}})}\|\partial^{\leq k}A\|_{L^{3}(C_{T_{1}})},$
@@
(4.17) $\|G_{I}\|_{L^{\frac{3}{2}}(C_{T_{1}})}\lesssim\|\psi\|_{L^{6}(C_{T_{1}})}\|\partial^{\leq k}\psi\|_{L^{2}(C_{T_{1}})}.$
@@
$\begin{aligned} \frac{1}{p_{1}}&=\frac{j}{3k}=\frac{j}{3}+\theta_{1}\Big{(}\frac{1}{3}-\frac{k}{3}\Big{)},\qquad&\theta_{1}&=\frac{j}{k},\\ \frac{1}{p_{2}}&=\frac{k-j}{3k}+\frac{1}{6}=\frac{k-j}{3}+\theta_{2}\Big{(}\frac{1}{2}-\frac{k}{3}\Big{)}+\frac{1-\theta_{2}}{6},\qquad&\theta_{2}&=\frac{k-j}{k}.\end{aligned}$
@@
$\|\partial^{I^{\prime}}A_{\mu}\partial^{I-I^{\prime}}\psi\|_{L^{2}(C_{T_{1}})}\leq$ $\ \|\partial^{I^{\prime}}A_{\mu}\|_{L^{p_{1}}(C_{T_{1}})}\|\partial^{I-I^{\prime}}\psi\|_{L^{p_{2}}(C_{T_{1}})}$
$\lesssim$ $\ \|\partial^{\leq k}A\|_{L^{3}}^{\theta_{1}}\|A\|_{L^{\infty}(C_{T_{1}})}^{1-\theta_{1}}\|\partial^{\leq k}\psi\|_{L^{2}(C_{T_{1}})}^{\theta_{2}}\|\psi\|_{L^{6}(C_{T_{1}})}^{1-\theta_{2}}$
$\lesssim$ $\ \|A\|_{L^{\infty}(C_{T_{1}})}\|\partial^{\leq k}\psi\|_{L^{2}(C_{T_{1}})}+\|\psi\|_{L^{6}(C_{T_{1}})}\|\partial^{\leq k}A\|_{L^{3}(C_{T_{1}})}$
@@
$\begin{aligned} \frac{1}{p_{1}}&=\frac{j}{3k}+\frac{1}{6}=\frac{j}{3}+\theta_{1}\Big{(}\frac{1}{2}-\frac{k}{3}\Big{)}+\frac{1-\theta_{1}}{6},\qquad&\theta_{1}&=\frac{j}{k},\\ \frac{1}{p_{2}}&=\frac{1}{2}-\frac{j}{3k}=\frac{k-j}{3}+\theta_{2}\Big{(}\frac{1}{2}-\frac{k}{3}\Big{)}+\frac{1-\theta_{2}}{6},&\theta_{2}&=\frac{k-j}{k},\end{aligned}$
@@
$\|\partial^{I^{\prime}}\overline{\psi}\gamma^{\mu}\partial^{I-I^{\prime}}\psi\|_{L^{\frac{3}{2}}(C_{T_{1}})}\lesssim$ $\ \|\partial^{I^{\prime}}\psi\|_{L^{p_{1}}(C_{T_{1}})}\|\partial^{I-I^{\prime}}\psi\|_{L^{p_{2}}(C_{T_{1}})}$
$\lesssim$ $\ \|\partial^{\leq k}\psi\|_{L^{2}(C_{T_{1}})}^{\theta_{1}}\|\psi\|_{L^{6}(C_{T_{1}})}^{1-\theta_{1}}\|\partial^{\leq k}\psi\|_{L^{2}(C_{T_{1}})}^{\theta_{2}}\|\psi\|_{L^{6}(C_{T_{1}})}^{1-\theta_{2}}$
$\lesssim$ $\ \|\partial^{\leq k}\psi\|_{L^{2}(C_{T_{1}})}\|\psi\|_{L^{6}(C_{T_{1}})},$
again as needed. This concludes the proof of (4.16) and (4.17) and thus the proof of the lemma.
The estimates including the time derivatives now follow inductively from (3.1). ∎
@@ 4.4. Bounds for vector fields
(4.18) $\|\Omega^{J}u\|_{L^{p}(C_{T}^{+})}\lesssim\|\Omega^{\leq m}u\|_{L^{r}(C_{T}^{+})}^{\theta}\|u\|_{L^{q}(C_{T}^{+})}^{1-\theta}.$
@@
(4.21) $T_{1}^{\frac{1}{2}}\|F_{J}\|_{L^{2}(C_{T_{1}})}\lesssim\varepsilon\,(\|\Omega^{\leq k}\psi\|_{L^{\infty}L^{2}(C_{<T_{1}})}+\|\Omega^{\leq k}A\|_{L^{\infty}L^{3}(C_{<T_{1}})}),$
(4.22) $T_{1}^{\frac{1}{3}}\|G_{J}\|_{L^{\frac{3}{2}}(C_{T_{1}})}\lesssim\varepsilon\,\|\Omega^{\leq k}\psi\|_{L^{\infty}L^{2}(C_{<T_{1}})}.$
@@
$\|\Omega^{\leq k}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}+\|\Omega^{\leq k}A\|_{L^{\infty}L^{3}(C_{T_{1}})}$ $\lesssim\|\Omega^{\leq k}\psi(T_{1})\|_{L^{2}}+\|\nabla_{t,x}\,\Omega^{\leq k}A(T_{1})\|_{\dot{H}^{-\frac{1}{2}}}$
$\qquad+\sum_{|J|\leq k}(\|F_{J}\|_{L^{1}L^{2}(C_{T_{1}})}+\|G_{J}\|_{L^{1}L^{\frac{3}{2}}(C_{T_{1}})}).$
@@
(4.23) $\|\Omega^{\leq k}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}+\|\Omega^{\leq k}A\|_{L^{\infty}L^{3}(C_{T_{1}})}\lesssim\|\Omega^{\leq k}\psi(T_{1})\|_{L^{2}}+\|\nabla_{t,x}\,\Omega^{\leq k}A(T_{1})\|_{\dot{H}^{-\frac{1}{2}}}.$
@@
$T_{1}^{\frac{1}{2}}\|F_{J}\|_{L^{2}(C_{T_{1}})}+T_{1}^{\frac{1}{3}}\|G_{J}\|_{L^{\frac{3}{2}}(C_{T_{1}})}\lesssim\varepsilon(\|\Omega^{\leq k}\psi(T_{1})\|_{L^{2}}+\|\nabla_{t,x}\,\Omega^{\leq k}A(T_{1})\|_{\dot{H}^{-\frac{1}{2}}}).$
@@
$\qquad+C\sum_{|J|\leq k}(\|F_{J}\|_{L^{1}L^{2}([T_{1},2T_{1}]\times{\mathbb{R}}^{3})}+\|G_{J}\|_{L^{1}L^{\frac{3}{2}}([T_{1},2T_{1}]\times{\mathbb{R}}^{3})})$
@@
$T_{1}^{\frac{1}{2}}\|\Omega^{J_{1}}A_{\mu}\gamma^{\mu}\Omega^{J-J_{1}}\psi\|_{L^{2}(C_{T_{1}})}\lesssim T_{1}^{\frac{1}{2}}\|\Omega^{J_{1}}A\|_{L^{p_{1}}(C_{T_{1}})}\|\Omega^{J-J_{1}}\psi\|_{L^{p_{2}}(C_{T_{1}})}.$
@@
$\|\Omega^{J_{1}}A\|_{L^{p_{1}}(C_{T_{1}})}$ $\lesssim\|\Omega^{\leq k}A\|_{L^{3}(C_{T_{1}})}^{\frac{j}{k}}\|A\|_{L^{\infty}(C_{T_{1}})}^{1-\frac{j}{k}}$
$\lesssim T_{1}^{\frac{1}{3}\frac{j}{k}}\|\Omega^{\leq k}A\|_{L^{\infty}L^{3}(C_{T_{1}})}^{\frac{j}{k}}(C_{0}\varepsilon T_{1}^{-1})^{1-\frac{j}{k}},$
@@
$\|\Omega^{J-J_{1}}\psi\|_{L^{p_{2}}(C_{T_{1}})}$ $\lesssim\|\Omega^{\leq k}\psi\|_{L^{2}(C_{T_{1}})}^{1-\frac{j}{k}}\|\psi\|_{L^{6}(C_{T_{1}})}^{\frac{j}{k}}$
$\lesssim T_{1}^{\frac{1}{2}(1-\frac{j}{k})}\|\Omega^{\leq k}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}^{1-\frac{j}{k}}T_{1}^{\frac{1}{6}\frac{j}{k}}(C_{0}\varepsilon T_{1}^{-1})^{\frac{j}{k}}.$
@@
$T_{1}^{\frac{1}{2}}\|\Omega^{J_{1}}A_{\mu}\gamma^{\mu}\Omega^{J-J_{1}}\psi\|_{L^{2}(C_{T_{1}})}$ $\lesssim C_{0}\varepsilon\|\Omega^{\leq k}A\|_{L^{\infty}L^{3}(C_{T_{1}})}^{\frac{j}{k}}\|\Omega^{\leq k}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}^{1-\frac{j}{k}}$
$\lesssim C_{0}\varepsilon(\|\Omega^{\leq k}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}+\|\Omega^{\leq k}A\|_{L^{\infty}L^{3}(C_{T_{1}})}).$
@@
$T_{1}^{\frac{1}{3}}\|\Omega^{J_{1}}\overline{\psi}\gamma_{\mu}\Omega^{J-J_{1}}\psi\|_{L^{\frac{3}{2}}(C_{T_{1}})}\lesssim T_{1}^{\frac{1}{3}}\|\Omega^{J_{1}}\psi\|_{L^{p_{1}}(C_{T_{1}})}\|\Omega^{J-J_{1}}\psi\|_{L^{p_{2}}(C_{T_{1}})}.$
@@
$\|\Omega^{J_{1}}\psi\|_{L^{p_{1}}(C_{T_{1}})}$ $\lesssim\|\Omega^{\leq k}\psi\|_{L^{2}(C_{T_{1}})}^{\frac{j}{k}}\|\psi\|_{L^{6}(C_{T_{1}})}^{1-\frac{j}{k}}$
$\lesssim T_{1}^{\frac{1}{2}\frac{j}{k}}\|\Omega^{\leq k}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}^{\frac{j}{k}}T_{1}^{\frac{1}{6}(1-\frac{j}{k})}(C_{0}\varepsilon T_{1}^{-1})^{1-\frac{j}{k}},$
@@
$\|\Omega^{J-J_{1}}\psi\|_{L^{p_{2}}(C_{T_{1}})}\lesssim T_{1}^{\frac{1}{2}(1-\frac{j}{k})}\|\Omega^{\leq k}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}^{1-\frac{j}{k}}T_{1}^{\frac{1}{6}\frac{j}{k}}(C_{0}\varepsilon T_{1}^{-1})^{\frac{j}{k}}.$
@@
$T_{1}^{\frac{1}{3}}\|\Omega^{J_{1}}\overline{\psi}\gamma_{\mu}\Omega^{J-J_{1}}\psi\|_{L^{\frac{3}{2}}(C_{T_{1}})}\lesssim C_{0}\varepsilon\|\Omega^{\leq k}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}.$

### v5

@@ 4.3. Bounds for derivatives
Next, using Gagliardo-Nirenberg estimates in space-time, we derive energy estimates for $(\partial^{I}\psi,\partial^{I}A)$.
@@
(4.13) $T_{1}^{\frac{1}{2}}\|F_{I}\|_{L^{2}_{t,x}(C_{T_{1}})}+T_{1}^{\frac{1}{3}}\|G_{I}\|_{L^{\frac{3}{2}}_{t,x}(C_{T_{1}})}\lesssim C_{0}\varepsilon\|\partial^{\leq k}(\psi,A)\|_{X_{T_{1}}},\qquad|I|\leq k.$
@@
$T_{1}^{\frac{1}{2}}\|F_{I}\|_{L^{2}_{t,x}(C_{T_{1}})}+T_{1}^{\frac{1}{3}}\|G_{I}\|_{L^{\frac{3}{2}}_{t,x}(C_{T_{1}})}\lesssim C_{0}\varepsilon\|\partial^{\leq k}(\psi,A)(T_{1})\|_{\mathscr{H}^{0}}.$
@@
This last inequality forms the basis for our discrete Gronwall argument. For any $t\in[1,T]$, we fix $l\in{\mathbb{N}}$ with $2^{l-1}\leq t<2^{l}$ and set $t_{1}=2^{-l}t\in[0,1)$. Iterating the above estimate, we infer
$\|\partial^{\leq k}(\psi,A)(t)\|_{\mathscr{H}^{0}}\leq(1+c\varepsilon)^{l}\|\partial^{\leq k}(\psi,A)(t_{1})\|_{\mathscr{H}^{0}}.$
Since $(1+c\varepsilon)^{l}\lesssim t^{c^{\prime}\varepsilon}$ with $c^{\prime}\approx c\approx C_{0}$ and $\|\partial^{\leq k}(\psi,A)\|_{L^{\infty}_{t}([0,1])\mathscr{H}^{0}}\lesssim\|\partial^{\leq k}(\psi,A)(0)\|_{\mathscr{H}^{0}}$ by local well-posedness, we arrive at
$\|\partial^{\leq k}(\psi,A)(t)\|_{\mathscr{H}^{0}}\lesssim\langle t\rangle^{c\varepsilon}\|\partial^{\leq k}(\psi,A)(0)\|_{\mathscr{H}^{0}},\qquad 0\leq t\leq T$
with a constant $c\approx C_{0}$. Together with (4.14), this implies the desired estimate (4.12).
@@
(4.16) $\|F_{I}\|_{L^{2}_{t,x}(C_{T_{1}})}\lesssim\|A\|_{L^{\infty}_{t,x}(C_{T_{1}})}\|\partial^{\leq k}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}+\|\psi\|_{L^{6}_{t,x}(C_{T_{1}})}\|\partial^{\leq k}A\|_{L^{3}_{t,x}(C_{T_{1}})},$
@@
(4.17) $\|G_{I}\|_{L^{\frac{3}{2}}_{t,x}(C_{T_{1}})}\lesssim\|\psi\|_{L^{6}_{t,x}(C_{T_{1}})}\|\partial^{\leq k}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}.$
@@
$\begin{aligned} \frac{1}{p_{1}}&=\frac{j}{3k}=\frac{j}{4}+\theta_{1}\Big(\frac{1}{3}-\frac{k}{4}\Big),\qquad&\theta_{1}&=\frac{j}{k},\\ \frac{1}{p_{2}}&=\frac{k-j}{3k}+\frac{1}{6}=\frac{k-j}{4}+\theta_{2}\Big(\frac{1}{2}-\frac{k}{4}\Big)+\frac{1-\theta_{2}}{6},\qquad&\theta_{2}&=\frac{k-j}{k}.\end{aligned}$
@@
$\|\partial^{I^{\prime}}A_{\mu}\partial^{I-I^{\prime}}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}\leq$ $\ \|\partial^{I^{\prime}}A_{\mu}\|_{L^{p_{1}}_{t,x}(C_{T_{1}})}\|\partial^{I-I^{\prime}}\psi\|_{L^{p_{2}}_{t,x}(C_{T_{1}})}$
$\lesssim$ $\ \|\partial^{\leq k}A\|_{L^{3}_{t,x}(C_{T_{1}})}^{\theta_{1}}\|A\|_{L^{\infty}_{t,x}(C_{T_{1}})}^{1-\theta_{1}}\|\partial^{\leq k}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}^{\theta_{2}}\|\psi\|_{L^{6}_{t,x}(C_{T_{1}})}^{1-\theta_{2}}$
$\lesssim$ $\ \|A\|_{L^{\infty}_{t,x}(C_{T_{1}})}\|\partial^{\leq k}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}+\|\psi\|_{L^{6}_{t,x}(C_{T_{1}})}\|\partial^{\leq k}A\|_{L^{3}_{t,x}(C_{T_{1}})}$
@@
$\begin{aligned} \frac{1}{p_{1}}&=\frac{j}{3k}+\frac{1}{6}=\frac{j}{4}+\theta_{1}\Big(\frac{1}{2}-\frac{k}{4}\Big)+\frac{1-\theta_{1}}{6},\qquad&\theta_{1}&=\frac{j}{k},\\ \frac{1}{p_{2}}&=\frac{1}{2}-\frac{j}{3k}=\frac{k-j}{4}+\theta_{2}\Big(\frac{1}{2}-\frac{k}{4}\Big)+\frac{1-\theta_{2}}{6},&\theta_{2}&=\frac{k-j}{k},\end{aligned}$
@@
$\|\partial^{I^{\prime}}\overline{\psi}\gamma^{\mu}\partial^{I-I^{\prime}}\psi\|_{L^{\frac{3}{2}}_{t,x}(C_{T_{1}})}\lesssim$ $\ \|\partial^{I^{\prime}}\psi\|_{L^{p_{1}}_{t,x}(C_{T_{1}})}\|\partial^{I-I^{\prime}}\psi\|_{L^{p_{2}}_{t,x}(C_{T_{1}})}$
$\lesssim$ $\ \|\partial^{\leq k}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}^{\theta_{1}}\|\psi\|_{L^{6}_{t,x}(C_{T_{1}})}^{1-\theta_{1}}\|\partial^{\leq k}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}^{\theta_{2}}\|\psi\|_{L^{6}_{t,x}(C_{T_{1}})}^{1-\theta_{2}}$
$\lesssim$ $\ \|\partial^{\leq k}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}\|\psi\|_{L^{6}_{t,x}(C_{T_{1}})},$
again as needed. This concludes the proof of (4.16) and (4.17) and thus the proof of the lemma. ∎
@@ 4.4. Bounds for vector fields
(4.18) $\|\Omega^{J}u\|_{L^{p}_{t,x}(C_{T}^{+})}\lesssim\|\Omega^{\leq m}u\|_{L^{r}_{t,x}(C_{T}^{+})}^{\theta}\|u\|_{L^{q}_{t,x}(C_{T}^{+})}^{1-\theta}.$
@@
(4.21) $T_{1}^{\frac{1}{2}}\|F_{J}\|_{L^{2}_{t,x}(C_{T_{1}})}\lesssim\varepsilon\,(\|\Omega^{\leq k}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{<T_{1}})}+\|\Omega^{\leq k}A\|_{L^{\infty}_{t}L^{3}_{x}(C_{<T_{1}})}),$
(4.22) $T_{1}^{\frac{1}{3}}\|G_{J}\|_{L^{\frac{3}{2}}_{t,x}(C_{T_{1}})}\lesssim\varepsilon\,\|\Omega^{\leq k}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{<T_{1}})}.$
@@
$\|\Omega^{\leq k}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}+\|\Omega^{\leq k}A\|_{L^{\infty}_{t}L^{3}_{x}(C_{T_{1}})}$ $\lesssim\|\Omega^{\leq k}\psi(T_{1})\|_{L^{2}}+\|\nabla_{t,x}\,\Omega^{\leq k}A(T_{1})\|_{\dot{H}^{-\frac{1}{2}}}$
$\qquad+\sum_{|J|\leq k}(\|F_{J}\|_{L^{1}_{t}L^{2}_{x}(C_{T_{1}})}+\|G_{J}\|_{L^{1}_{t}L^{\frac{3}{2}}_{x}(C_{T_{1}})}).$
@@
(4.23) $\|\Omega^{\leq k}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}+\|\Omega^{\leq k}A\|_{L^{\infty}_{t}L^{3}_{x}(C_{T_{1}})}\lesssim\|\Omega^{\leq k}\psi(T_{1})\|_{L^{2}}+\|\nabla_{t,x}\,\Omega^{\leq k}A(T_{1})\|_{\dot{H}^{-\frac{1}{2}}}.$
@@
$T_{1}^{\frac{1}{2}}\|F_{J}\|_{L^{2}_{t,x}(C_{T_{1}})}+T_{1}^{\frac{1}{3}}\|G_{J}\|_{L^{\frac{3}{2}}_{t,x}(C_{T_{1}})}\lesssim\varepsilon(\|\Omega^{\leq k}\psi(T_{1})\|_{L^{2}}+\|\nabla_{t,x}\,\Omega^{\leq k}A(T_{1})\|_{\dot{H}^{-\frac{1}{2}}}).$
@@
$\qquad+C\sum_{|J|\leq k}(\|F_{J}\|_{L^{1}_{t}L^{2}_{x}([T_{1},2T_{1}]\times{\mathbb{R}}^{3})}+\|G_{J}\|_{L^{1}_{t}L^{\frac{3}{2}}_{x}([T_{1},2T_{1}]\times{\mathbb{R}}^{3})})$
@@
$T_{1}^{\frac{1}{2}}\|\Omega^{J_{1}}A_{\mu}\gamma^{\mu}\Omega^{J-J_{1}}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}\lesssim T_{1}^{\frac{1}{2}}\|\Omega^{J_{1}}A\|_{L^{p_{1}}_{t,x}(C_{T_{1}})}\|\Omega^{J-J_{1}}\psi\|_{L^{p_{2}}_{t,x}(C_{T_{1}})}.$
@@
$\|\Omega^{J_{1}}A\|_{L^{p_{1}}_{t,x}(C_{T_{1}})}$ $\lesssim\|\Omega^{\leq k}A\|_{L^{3}_{t,x}(C_{T_{1}})}^{\frac{j}{k}}\|A\|_{L^{\infty}_{t,x}(C_{T_{1}})}^{1-\frac{j}{k}}$
$\lesssim T_{1}^{\frac{1}{3}\frac{j}{k}}\|\Omega^{\leq k}A\|_{L^{\infty}_{t}L^{3}_{x}(C_{T_{1}})}^{\frac{j}{k}}(C_{0}\varepsilon T_{1}^{-1})^{1-\frac{j}{k}},$
@@
$\|\Omega^{J-J_{1}}\psi\|_{L^{p_{2}}_{t,x}(C_{T_{1}})}$ $\lesssim\|\Omega^{\leq k}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}^{1-\frac{j}{k}}\|\psi\|_{L^{6}_{t,x}(C_{T_{1}})}^{\frac{j}{k}}$
$\lesssim T_{1}^{\frac{1}{2}(1-\frac{j}{k})}\|\Omega^{\leq k}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}^{1-\frac{j}{k}}T_{1}^{\frac{1}{6}\frac{j}{k}}(C_{0}\varepsilon T_{1}^{-1})^{\frac{j}{k}}.$
@@
$T_{1}^{\frac{1}{2}}\|\Omega^{J_{1}}A_{\mu}\gamma^{\mu}\Omega^{J-J_{1}}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}$ $\lesssim C_{0}\varepsilon\|\Omega^{\leq k}A\|_{L^{\infty}_{t}L^{3}_{x}(C_{T_{1}})}^{\frac{j}{k}}\|\Omega^{\leq k}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}^{1-\frac{j}{k}}$
$\lesssim C_{0}\varepsilon(\|\Omega^{\leq k}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}+\|\Omega^{\leq k}A\|_{L^{\infty}_{t}L^{3}_{x}(C_{T_{1}})}).$
@@
$T_{1}^{\frac{1}{3}}\|\Omega^{J_{1}}\overline{\psi}\gamma_{\mu}\Omega^{J-J_{1}}\psi\|_{L^{\frac{3}{2}}_{t,x}(C_{T_{1}})}\lesssim T_{1}^{\frac{1}{3}}\|\Omega^{J_{1}}\psi\|_{L^{p_{1}}_{t,x}(C_{T_{1}})}\|\Omega^{J-J_{1}}\psi\|_{L^{p_{2}}_{t,x}(C_{T_{1}})}.$
@@
$\|\Omega^{J_{1}}\psi\|_{L^{p_{1}}_{t,x}(C_{T_{1}})}$ $\lesssim\|\Omega^{\leq k}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}^{\frac{j}{k}}\|\psi\|_{L^{6}_{t,x}(C_{T_{1}})}^{1-\frac{j}{k}}$
$\lesssim T_{1}^{\frac{1}{2}\frac{j}{k}}\|\Omega^{\leq k}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}^{\frac{j}{k}}T_{1}^{\frac{1}{6}(1-\frac{j}{k})}(C_{0}\varepsilon T_{1}^{-1})^{1-\frac{j}{k}},$
@@
$\|\Omega^{J-J_{1}}\psi\|_{L^{p_{2}}_{t,x}(C_{T_{1}})}\lesssim T_{1}^{\frac{1}{2}(1-\frac{j}{k})}\|\Omega^{\leq k}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}^{1-\frac{j}{k}}T_{1}^{\frac{1}{6}\frac{j}{k}}(C_{0}\varepsilon T_{1}^{-1})^{\frac{j}{k}}.$
@@
$T_{1}^{\frac{1}{3}}\|\Omega^{J_{1}}\overline{\psi}\gamma_{\mu}\Omega^{J-J_{1}}\psi\|_{L^{\frac{3}{2}}_{t,x}(C_{T_{1}})}\lesssim C_{0}\varepsilon\|\Omega^{\leq k}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}.$